OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 5 interactions. If there are 4 children, there can be 14 interactions. Which recursive equation represents the pattern?

OpenStudy (anonymous):

These are the options an = an – 1 + 2(n – 1) an = an – 1 + (n – 1) 2 an = an – 1 + 2(n – 1) an = an – 1 + (2n – 1)

ganeshie8 (ganeshie8):

plugin "an - 1" and "n" in the options and see which one matches. an = an – 1 + 2(n – 1) doesnt match an = an – 1 + (n – 1) 2 matches